On the front row of the concert the sound level reaches a value of 110 dB. In the back row thw sound intensity is reduced by a factor of 1000. What is the decibel level on the back row?
1
Expert's answer
2011-06-20T12:42:29-0400
Decibel level defined as

L1 = 10 log10(I1/I0) dB

where I1 and I0 are the intensities.
I0 is the standard reference sound intensity

L1 = 10*log10 (I1/I0) = 110 dB - sound level on the front row.
L2 = 10*log10 (I2/I0) = 110 dB - sound level on the back row.
I2=I1/1000
So
L2=10*log10 (I1/1000)/I0 = 10*log10(I1/I0) + 10*log10(1/1000) = 110 - 3 = 107 dB
Because
Answer : 107 dB.